### v1

# Finite-blocklength Fluid Antenna Systems With Spatial Block-Correlation Channel Model

Abstract
Massive connectivity with ultra-low latency and high reliability necessitates fundamental advances in future communication networks operating under finite-blocklength (FBL) transmission. Fluid antenna systems (FAS) have emerged as a promising enabler, offering superior spectrum and energy efficiency in short-packet/FBL scenarios. In this work, by leveraging the simplicity and accuracy of block-correlation channel modeling, we rigorously bound the performance limits of FBL-FAS from a statistical perspective, focusing on two key performance metrics: block error rate (BLER) and outage probability (OP). Furthermore, we introduce a novel complex-integral simplification method based on Gauss-Laguerre quadrature, which achieves higher approximation accuracy compared to existing Taylor-expansion-based approaches. Numerical results validate the robustness of the proposed analysis and clearly demonstrate the superiority of FBL-FAS over conventional multiple-antenna systems with fixed antenna placement.
I Introduction
Fluid antenna systems (FAS) represent a communication architecture that achieves substantial spatial diversity gains by dynamically altering radiation characteristics through flexible antenna structures and adaptive protocol designs [1] . Unlike conventional approaches that seek to avoid antenna correlation in compact spaces, FAS intentionally embraces and exploits it to yield significant channel gains [2], thereby enabling robust service quality with reduced hardware overhead and simplified protocol design. Recent work [3] investigated the performance limits of FAS under the finite blocklength (FBL) regime, demonstrating that fluid antennas can simultaneously secure high spectrum and power efficiency. Furthermore, FBL theories [4, 5] provide both fundamental insights and practical guidelines for system design to meet the requirements of massive access, ultra-low latency, and high reliability. Unlike prior studies that primarily adopt outage probability (OP) as the performance metric for FAS, [3] introduces block error rate (BLER) as a universal and analytically tractable metric with a closed-form expression. The main scope and contributions of this work are summarized as follows:
    In this work, we focus on the FBL-FAS problem [3] under the block-correlation channel model [6] of FAS. With accurate channel modeling, the performance limits of FBL-FAS are statistically characterized by providing tighter upper bounds without requiring prior channel state information (CSI).
    We derive the PDF expression of the channel response after antenna/port selection, which is concisely represented by a non-central chi-square distribution.
    Furthermore, we introduce a more general integral simplification method inspired by the Gauss-Laguerre approach. Compared with the previously used Taylor-expansion-assisted method [3], our method avoids amplified approximation errors from high-order expansion terms under large antenna correlation, and achieves improved approximation accuracy.
The remainder of the paper is organized as follows: In Section II, FBL-FASs under block-correlation channel modeling are explained. Subsequently, our mains results are introduced in Section III while explicit derivations are deferred to Section VI. Numerical results are illustrated and explained in Section IV and the conclusion is drawn in Section V.
II System Descriptions
In this work, we consider an uplink transmission scenario where a base station (BS) equipped with a conventional single antenna serves $U$ single-antenna user equipments with blocklength $M$. Each UE employs a fluid antenna with $N$ preset configurable locations/ports that are evenly distributed along a linear array of length $W$, where $W$ denotes the antenna length normalized by the wavelength at the operating frequency. Each port is regarded as an ideal antenna; however, due to limited hardware resources such as the radio frequency (RF) chain, only one port will be activated at a time.
II-A Spatial Block-Correlation Channel Model for Single Antenna $N$-Ports FAS
Our starting point builds upon the FBL-FAS analytical model in [3, 2]. In contrast, we extend the analysis to incorporate the spatial block-correlation channel model in [6]. Since this channel model differs from that in [3], it is necessary to illustrate how the block-correlation model can be adopted to generate random channel realizations within the FBL-FAS framework. Let $\boldsymbol{g}_{u}\in\mathbb{C}^{N}$ denote the channel coefficient vector of the $u$-th user, and let $\boldsymbol{\Sigma}\in\mathbb{C}^{N\times N}$ represent the corresponding spatial correlation matrix. Owing to the even placement of ports, the correlation matrix $\boldsymbol{\Sigma}$ takes the form of a Toeplitz matrix:
$\boldsymbol{\Sigma}=\begin{pmatrix}a(0)&a(1)&a(2)&...&a(N-1)\\ a(-1)&a(0)&a(1)&...&a(N-2)\\ \vdots&\ddots&&\vdots\\ a(-N+1)&a(-N+2)&...&a(-1)&a(0)\end{pmatrix}$, (1)
where the generation function is:
$a(n)=\operatorname{sinc}\left(\frac{2\pi nW}{N-1}\right).$ (2)
With correlation matrix $\boldsymbol{\Sigma}$, one could generate $\boldsymbol{g}_{u}$ feasibly via eigenvalue-based construction [1, Eq. 5]:
$\boldsymbol{g}_{u}=\boldsymbol{Q}\boldsymbol{\Lambda}^{\frac{1}{2}}\boldsymbol{g}_{0},$ (3)
Here, $\boldsymbol{Q}$ denotes the matrix of left eigenvectors obtained from the eigenvalue decomposition of the correlation matrix $\boldsymbol{\Sigma}=\boldsymbol{Q}\boldsymbol{\Lambda}\boldsymbol{Q}^{\mathrm{H}}$, and $\boldsymbol{g}_{0}\in\mathbb{C}^{N}\sim\mathcal{CN}\left(\boldsymbol{0},\sigma^{2}\boldsymbol{I}\right)$. However, directly analyzing the statistical distribution based on $\boldsymbol{\Sigma}$ is infeasible and overly complicated for any tractable analyses.
Fortunately, the principal components of $\boldsymbol{\Sigma}$ can be captured by only a small number of dominant eigenvalues and their corresponding eigenvectors [6, Fig. 3], [1]. Hence, the correlation matrix $\boldsymbol{\Sigma}$ can be approximated by a block-diagonal matrix:
$\widehat{\boldsymbol{\Sigma}}\in\mathbb{R}^{N\times N}=\begin{pmatrix}\mathbf{A}_{1}&\mathbf{0}&\cdots&\mathbf{0}\\ \mathbf{0}&\mathbf{A}_{2}&\cdots&\mathbf{0}\\ \vdots&&\ddots&\vdots\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{A}_{B}\end{pmatrix}$, (4)
Each diagonal block $\boldsymbol{A}_{b},~b\in\{1,2,\ldots,B\}$, is a constant correlation matrix of size $L_{b}\times L_{b}$, satisfying $\sum_{b=1}^{B}L_{b}=N$ (omit rounding errors). Intuitively, the correlation matrix characterizes the correlation among neighboring ports, where the correlation is typically strong due to the compact port spacing. The matrix $\widehat{\boldsymbol{\Sigma}}$ represents the fitted approximation of $\boldsymbol{\Sigma}$ according to [6, Eq. 20], while offering a more convenient structure for analytical treatment. In particular, the block-correlation matrix exhibits the following common form:
$\mathbf{A}_{b}=\begin{pmatrix}1&\mu_{b}^{2}&\cdots&\mu_{b}^{2}\\ \mu_{b}^{2}&1&\cdots&\mu_{b}^{2}\\ \vdots&&\ddots&\vdots\\ \mu_{b}^{2}&\cdots&\mu_{b}^{2}&1\end{pmatrix}$, (5)
where $\mu_{b}^{2}$ is the correlation constant. Two important assumptions adopted throughout the remainder of the paper are summarized as follows:
    Without loss of generality, $\mu_{b}^{2},~b\in\{1,2,\ldots\}$ are assumed identical, i.e., $\mu_{1}^{2}=\mu_{2}^{2}=\cdots=\mu^{2}\in\left(0.95,0.99\right)$, as established in [6, Sec. III-B-2)].
    The block size $L_{b}$ is adaptive to the desired approximation accuracy, which follows the procedure in [6, Algorithm 1]. For implementation details, the reader is kindly referred to the open-source code provided therein.
II-B Finite-Blocklength Single Antenna FAS Signal Model
Assuming synchronous transmission, $U$ single-antenna users transmit $M$-length signals, and the received signal can be expressed as
$\boldsymbol{y}=g_{i,k}\boldsymbol{x}_{i}+\sum_{u\neq i}^{U}g_{u,k}\boldsymbol{x}_{u}+\boldsymbol{\eta},$ (6)
where $\boldsymbol{x}_{u}$ denotes the Gaussian codeword of user $u$ with zero mean and variance $\sigma_{c}^{2}=\tfrac{1}{M}$, such that $\|\boldsymbol{x}_{u}\|^{2}_{2}=1$. The term $\boldsymbol{\eta}$ represents the additive white Gaussian noise (AWGN) with zero mean and variance $\tfrac{\sigma^{2}_{\eta}}{M}$, i.e., $\|\boldsymbol{\eta}\|^{2}_{2}=\sigma^{2}_{\eta}$. With a slight abuse of notation, $g_{u,k}$ denotes the $k$-th channel coefficient of user $u$, where $k\in\{1,2,\ldots,N\}$. The channel coefficients are Rayleigh distributed with variance $\sigma^{2}$ and correlated by (4). Moreover, the signal-to-noise ratio (SNR) is defined as
$\mathrm{SNR}=\frac{\mathbb{E}\left\{\|g_{i,k}\boldsymbol{x}_{i}\|^{2}_{2}\right\}}{\mathbb{E}\left\{\|\boldsymbol{\eta}\|^{2}_{2}\right\}}=\frac{\sigma^{2}}{\sigma^{2}_{\eta}}.$ (7)
$f_{|g_{\mathrm{FAS}}|^{2}}(t)$ (8)
$=\sum_{b=1}^{B}\left\{\int_{0}^{\infty}\frac{e^{-\frac{r_{b}}{2}}L_{b}}{2}\left[F_{\chi_{2}^{\prime 2}}(\frac{t}{1-\mu^{2}};\frac{\mu^{2}r_{b}}{1-\mu^{2}})\right]^{L_{b}-1}\frac{f_{\chi_{2}^{\prime 2}}(\frac{t}{1-\mu^{2}};\frac{\mu^{2}r_{b}}{1-\mu^{2}})}{1-\mu^{2}}\mathrm{d}r_{b}\right\}\times\prod_{j\neq b}^{B}\left\{\int_{0}^{\infty}\frac{e^{-\frac{r_{j}}{2}}}{2}\left[F_{\chi_{2}^{\prime 2}}\left(\frac{t}{1-\mu^{2}};\frac{\mu^{2}r_{j}}{1-\mu^{2}}\right)\right]^{L_{j}}\mathrm{d}r_{j}\right\},$
III Main Results
In this section, we derive the PDF of the random variable $|g_{\mathrm{FAS}}|^{2}=\max\left\{|g_{u,1}|^{2},|g_{u,2}|^{2},\ldots,|g_{u,N}|^{2}\right\}$, based on which both the conditional BLER (with known channel response) and the statistical BLER (without prior CSI) are obtained. Furthermore, we investigate the outage OP expressions of FBL-FAS under the block-correlation channel model. In addition, we clarify why the integral simplification method in [3, Eq. 52], i.e., the Taylor-expansion-assisted mean value theorem for integrals, is not applicable to the block-correlation model. Instead, we propose a tailored simplification approach for block correlation model based on Gauss–Laguerre quadrature.
Figure 1: Illustration of block-correlation model’s CDF and the PDF (exact CDF in (13), derived PDF in (8) and Monte-Carlo samples.) with $N=10,~\mu^{2}=0.97,~W=0.5$.
III-A PDF of $|g_{\mathrm{FAS}}|^{2}$ and BLER under Block-Correlation Model
Result 1: The PDF of the random variable $|g_{\mathrm{FAS}}|^{2}=\max\left\{|g_{u,1}|^{2},|g_{u,2}|^{2},\ldots,|g_{u,N}|^{2}\right\}$ for FBL-FAS under the block-correlation channel model is given in (8), where $f_{\chi_{2}^{\prime 2}}(x;\lambda)$ and $F_{\chi_{2}^{\prime 2}}(x;\lambda)$ denote the PDF and CDF of a non-central chi-square distribution with $2$ degrees of freedom and non-centrality parameter $\lambda$, respectively.
Proof: See Appendix VI-A. $\blacksquare$
For illustration, the exact CDF in (13), the derived PDF in (8), and the Monte Carlo samples obtained via (3) are plotted in Fig. 1, where the channel model setups are generated with $N=10$, $\mu^{2}=0.97$, and $W=0.5$. Notably, consistent with [6], the CDF in (13) assumes $\sigma^{2}=2$, which will be adopted throughout the simulations in this work [^1] It is straightforward to extend the CDF and PDF to different variances $\sigma^{2}$, which can be done by a standard change of variables in the PDF/CDF..
Remark 1 According to [3, Theorem 5], for FBL-FAS with $U$ codewords to be detected under transmision of $M$ blocklength, the BLER can be upperbounded by the conditional (known channel response) error probability denoted by $\epsilon$:
$\epsilon\leq\sum_{{U}^{\prime}=0}^{U}\frac{{U}^{\prime}}{U}e^{{L}^{\prime}-M\log\left(1+\frac{0.25\sigma_{\eta^{\prime}}^{2}}{\sigma_{\eta}^{2}}\right)},$ (9)
where ${L}^{\prime}=\log\binom{U}{{U}^{\prime}}^{2}=\sum_{i=1}^{{U}^{\prime}-1}\log\frac{U-i}{{U^{\prime}-i}}$ and $\sigma^{2}_{\eta^{\prime}}=2{U}^{\prime}\sigma_{c}^{2}|g_{\mathrm{FAS}}|^{2}$. With the conditional BLER in (9), one could easily obtain the statistical BLER (without CSI) denoted by $\bar{\epsilon}$ below:
$\bar{\epsilon}=\sum_{{U}^{\prime}=0}^{U}\frac{{U}^{\prime}}{U}\int_{0}^{\infty}f_{|g_{\mathrm{FAS}}|^{2}}(t)e^{{L}^{\prime}-M\log\left(1+\frac{0.25\sigma_{\eta^{\prime}}^{2}}{\sigma_{\eta}^{2}(t)}\right)}\mathrm{d}t,$ (10)
where $f_{|g_{\mathrm{FAS}}|^{2}}(t)$ is the PDF of $|g_{\mathrm{FAS}}|^{2}$ and $\sigma_{\eta}^{2}(t)=2{U}^{\prime}\sigma_{c}^{2}t$.
III-B SINR-based Outage Probability and Integral Simplification
Remark 2 Let $\Gamma$ denote the SINR under FBL-FAS and $\gamma_{th}$ as the OP threshold, in accordance with [3, Appendix H], and let the OP event be defined as::
$\left\{\Gamma\leq\gamma_{\mathrm{th}}\right\}$ (11)
$=\left\{|g_{\mathrm{FAS}}|^{2}\leq\frac{\sigma_{\eta}^{2}\gamma_{\mathrm{th}}}{1-\left(U-1\right)\left(U\bar{\rho}+1\right)\gamma_{\mathrm{th}}}\right\},$
where $\bar{\rho}=\sqrt{\frac{\pi}{4M}}$ is the average codeword correlation (please refer to [3, Theorem 3] for details). Hereby, with CDF in (13), the OP for FBL-FAS under block-correlation channel model is calculated as:
$p^{\mathrm{FBL-FAS}}_{\mathrm{out}}$ (12)
$=\prod_{b=1}^{B}\int_{0}^{\infty}\frac{e^{-\frac{r_{b}}{2}}}{2}\times\left[1-Q_{1}\left(\sqrt{\frac{\mu^{2}r_{b}}{1-\mu^{2}}},\sqrt{\frac{t_{th}}{1-\mu^{2}}}\right)\right]^{L_{b}}\mathrm{d}r_{b},$
where $t_{th}=\frac{\sigma_{\eta}^{2}\gamma_{\mathrm{th}}}{1-\left(U-1\right)\left(U\bar{\rho}+1\right)\gamma_{\mathrm{th}}}$. Still, the expression in (12) has an infeasible product-integral structure. We next find an effective method to simplify the calculation.
Result 2: For the integral in (12), the simplification method originally adopted in [3, Theorem 7]—namely, the Taylor-expansion-assisted mean value theorem for integrals—is not applicable to the block-correlation model. In this work, we propose a novel approach for integral simplification based on Gauss–Laguerre quadrature, tailored to the analysis of FBL-FAS under the block-correlation model.
Details: See Appendix VI-B. $\blacksquare$
Figure 2: Illustrations of BLER performance:
-Different Lines: Red solid lines denote the statistical BLER of FBL-FAS, red markers represent the corresponding conditional BLER, and black dotted lines indicate the conditional BLER of the conventional $L$-antenna system.
-Different Figures: 1) BLER vs $U$: $W=1$, $M=5$, $N\in\{5,50\}$, $L\in\{1,2\}$ , and $\mathrm{SNR}=20$ dB; 2) BLER vs SNR: $W=0.5$, $M=5$, $U=10$, $N\in\{5,50,1000\}$, and $L\in\{1,2\}$; 3) BLER vs $N$: $W=1$, $M=5$, $U=10$, $L\in\{1,2\}$, and $\mathrm{SNR}=12$ dB. 4) BLER vs $W$: $M=5$, $U=10$, $N=5000$, $L\in\{1,2\}$, and $\mathrm{SNR}=12$ dB.
Figure 3: Illustration of OP performance:
-Different Lines: Red solid lines denote the OP of FBL-FAS, while black dotted lines indicate the OP of the MRC system.
-Different Figures: Up: OP versus SNR (dB) with $W=0.5$, $N\in\{5,50,500\}$, $L\in\{1,3,5\}$, $U=20$, $M=5$, and $\gamma_{th}=10^{-3}$; Down: OP versus the number of codewords $U$ with $W=0.5$, $N\in\{5,50,500,1000\}$, $L\in\{1,3,5\}$ (conventional antenna system), $\mathrm{SNR}=-35$ dB, $M=5$, and $\gamma_{th}=10^{-4}$.
IV Numerical Results
In this part, we illustrate the BLER and OP performance of FBL-FAS under the block-correlation model. The OP and BLER for a conventional $L$-antenna scheme using maximum ratio combining (MRC) [3, Section III-D] are included as benchmarks. For the Monte Carlo simulations, random channel realizations are generated according to (3). Unless otherwise specified, the default setup assumes channel variance $\sigma^{2}=2$ (consistent with [6]) and correlation parameter $\mu^{2}=0.97$.
IV-A BLER and OP Performance
The BLER and OP performance results are illustrated in Fig. 2 and Fig. 3, respectively, with parameter setups and explanations on curve and marker types in the caption part:
In Fig. 2, BLER results are presented under different numbers of codewords $U$, SNR levels, and numbers of ports $N$. Several common findings can be summarized: 1) The BLER can be significantly reduced with denser port placement and wider antenna length, i.e., increasing $N$ or $W$ improves the BLER performance, which can easily surpass that of conventional single-antenna systems and even multi-antenna systems. 2) With a sufficiently large number of ports $N$, the block-correlation channel model provides accurate predictions of BLER performance through the statistical expression in (10), thereby validating both the proposed approach and the accuracy of the block-correlation channel model.
In Fig. 3, the OP performance is illustrated under different SNR values and numbers of active codewords $U$. The key observations are as follows: 1) The single-antenna MRC system experiences an error floor due to user interference in finite-blocklength transmission. 2) FBL-FAS exhibits a much steeper OP decay as the SNR increases. Notably, even single-antenna FBL-FAS can achieve lower OP than a 5-antenna MRC system when $50$ or $500$ ports are available. 3) Compared with the 1-antenna MRC system, FBL-FAS achieves $10^{7}$ to $10^{4}$ times lower OP, and it even surpasses the 3-antenna MRC system with approximately 10 times better OP performance.
IV-B Integral Approximtion
To investigate the impact of $\mu$ on the approximation accuracy of different methods, we examine the integral component
$\int_{0}^{\infty}\frac{e^{-\frac{r_{b}}{2}}}{2}\left[1-Q_{1}\!\left(\sqrt{\tfrac{\mu^{2}r_{b}}{1-\mu^{2}}},\sqrt{\tfrac{t}{1-\mu^{2}}}\right)\right]^{L_{b}}\mathrm{d}r_{b}$
from (13) with $L_{b}=3$ and quadrature expansion order of $32$, as illustrated in Fig. 4. It can be observed that the Taylor expansion-assisted method becomes inaccurate as $\mu$ increases from $0.1$ to $0.97$. In contrast, the proposed Gauss–Laguerre-based approach remains effective across the entire range of correlation parameters.
Figure 4: Illustration of integral approximation under different correlation values in the block-correlation model, using the Taylor expansion-assisted method [3, Theorem 7] and the proposed Gauss–Laguerre method. The figures in the first row show the integral approximation while the figures in the second row show the corresponding approximation error, $|\tilde{f}(t)-f(t)|^{2}$. The $x$-axis denotes the range of $t$.
V Conclusion
In this work, the performance limits of FBL-FAS [3] are tightly bounded from a statistical perspective by considering the block-correlation channel model in [6]. The PDF expression of the channel gain after port selection is derived and concisely characterized by a non-central chi-square distribution, whose evaluation is further simplified using the Gauss-Laguerre approach to yield accurate approximations. Numerical results validate the tightness of the statistical bounds and highlight the significant spectrum and energy efficiency achieved by FASs in the FBL regime. Compared with conventional $L$-antenna systems, FBL-FAS emerges as a promising enabler for future wireless networks.
VI Appendix
VI-A Proof of Results 1
Remark A-1: Readily converted from [6, Eq. 43], the CDF of random variable $|g_{\mathrm{FAS}}|^{2}=\max\left\{|g_{u,1}|^{2},|g_{u,2}|^{2},\ldots,|g_{u,N}|^{2}\right\}$ is:
$\begin{aligned} &F_{|g_{\mathrm{FAS}}|^{2}}=P\left\{|g_{\mathrm{FAS}}|^{2}\leq t\right\}\\ &=\prod_{b=1}^{B}\int_{0}^{\infty}\frac{e^{-\frac{r_{b}}{2}}}{2}\times\left[1-Q_{1}\left(\sqrt{\frac{\mu^{2}r_{b}}{1-\mu^{2}}},\sqrt{\frac{t}{1-\mu^{2}}}\right)\right]^{L_{b}}\mathrm{d}r_{b}\end{aligned}.$ (13)
Subsequently, we derive the corresponding PDF by Remark A-1. For ease of descriptions, we denote (13) by:
$F_{|g_{\mathrm{FAS}}|}=\prod_{b=1}^{B}G_{b}(t),$ (14a)
$G_{b}(t)=\int_{0}^{\infty}\frac{1}{2}e^{-r_{b}/2}\left[1-Q_{1}\left(a_{b}(r_{b}),\sqrt{a_{t}}\right)\right]^{L_{b}}\mathrm{d}r_{b},$ (14b)
$a_{b}(r_{b})=\sqrt{\frac{\mu^{2}}{1-\mu^{2}}r_{b}},~\sqrt{a_{t}}=\sqrt{\frac{t}{1-\mu^{2}}}.$ (14c)
It is clear that by differentiating CDF, one could obtain the corresponding PDF by $f_{|g_{\mathrm{FAS}}|^{2}}=\frac{\mathrm{d}F_{|g_{\mathrm{FAS}}|^{2}}}{\mathrm{d}t}$. However, to simplify the derivative procedures, the following identities are considered:
    The equivalence between Marcum-$Q$ and non-central chi-square:
    $1-Q_{1}(a,\sqrt{x})=F_{\chi_{2}^{\prime 2}}(x;a^{2}),\quad x\geq 0,$ (15)
    where $F_{\chi_{2}^{\prime 2}}(x;a^{2})$ is the non-central chi-square CDF with degrees-of-freedom 2 and non-centrality parameter $a^{2}$.
    If $x$ in (15) is a function of variable $t$, i.e., $x(t)$, the following derivative can be established:
    $\frac{\mathrm{d}F_{\chi_{2}^{\prime 2}}(x(t);a^{2})}{\mathrm{d}t}=\frac{\mathrm{d}x(t)}{\mathrm{d}t}f_{\chi_{2}^{\prime 2}}(x(t);a^{2}),$ (16)
    where $f_{\chi_{2}^{\prime 2}}(x;a)=\frac{1}{2}e^{-\frac{x+a}{2}}I_{0}(\sqrt{ax})$ is the corresponding chi-square distribution’s PDF.
Thereby, if one consider indentity of (15) on (14a) and then use the derivative rule $f_{|g_{\mathrm{FAS}}|^{2}}(t)\equiv\frac{\mathrm{d}F_{|g_{\mathrm{FAS}}|^{2}}(t)}{\mathrm{d}t}=\sum_{b=1}^{B}\left(\frac{\mathrm{d}G_{b}(t)}{\mathrm{d}t}\right)\prod_{\begin{array}[]{c}j\neq b\end{array}}^{B}G_{j}(t)$ with identity in (16), (8) is obtained, which completes the proof on Result 1 in Section III-A.
VI-B Explanations on Result 2
Different from the constant channel correlation model in [3], the block-approximation model assumes the channel correlation to be a constant close to $1$, which aims to capture the correlation among ports in close vicinity. However, when $\mu^{2}\rightarrow 1$, if one continues to adopt the “frozen” point integral simplification method in [3, Theorem 7], the approximation becomes inaccurate due to the limitations of the Taylor expansion in [3, Eq. (56)] for $\mu^{2}\rightarrow 1$. Specifically, for the function component $1-Q_{1}\!\left(\sqrt{\tfrac{\mu^{2}r_{b}}{1-\mu^{2}}},\sqrt{\tfrac{t}{1-\mu^{2}}}\right)$ the constant terms $\tfrac{\mu^{2}}{1-\mu^{2}}\rightarrow+\infty$ or $\tfrac{1}{1-\mu^{2}}\rightarrow+\infty$ will excessively amplify the higher-order error terms after Taylor expansion, leading to failure of the approximation.
Therefore, it is both useful and interesting to explore ways to compensate for the drawbacks of Taylor-expansion-based methods. In this regard, we introduce how to simplify the integral calculation of expressions with structures similar to (12), inspired by the Gauss–Laguerre quadrature method [7, 8]:
    For an integrand of the form $e^{-x}f(x)$, the integral can be approximated as
    $\int_{0}^{\infty}e^{-x}f(x)\,\mathrm{d}x\;\approx\;\sum_{i=1}^{N_{GL}}w_{i}f(x_{i}),$ (17)
    where $x_{i}$ are the nodes and $w_{i}$ are the corresponding weights. The parameter $N_{GL}$ is the expansion order, which determines the approximation accuracy (typically, choosing $N_{GL}$ up to $64$ is sufficient).
    According to the Golub–Welsch theorem [7, 8], the values of $(x_{i},w_{i})$ can be derived by performing an eigenvalue decomposition on the following Jacobi matrix:
    $\boldsymbol{J}=\begin{pmatrix}\alpha_{0}&\sqrt{\beta_{1}}&0&\cdots&0\\ \sqrt{\beta_{1}}&\alpha_{1}&\sqrt{\beta_{2}}&\cdots&0\\ 0&\sqrt{\beta_{2}}&\alpha_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\ddots&\sqrt{\beta_{N-1}}\\ 0&0&\cdots&\sqrt{\beta_{N-1}}&\alpha_{N-1}\end{pmatrix}$, (18)
    where $\alpha_{n}=2n+1$ and $\beta_{n}=n^{2}$.
    The eigenvalue decomposition of (18), i.e., $\boldsymbol{J}=\boldsymbol{W}\boldsymbol{D}\boldsymbol{W}^{\mathrm{T}}$, provides the nodes $x_{i}$ as the $N$ largest eigenvalues (diagonal entries of $\boldsymbol{D}$). The corresponding weights $w_{i}$ are obtained from the squared elements of the first row of $\boldsymbol{W}$.

### v2

@@
 Abstract
-Massive connectivity with ultra-low latency and high reliability necessitates fundamental advances in future communication networks operating under finite-blocklength (FBL) transmission.
-Fluid antenna systems (FAS) have emerged as a promising enabler, offering superior spectrum and energy efficiency in short-packet/FBL scenarios.
-In this work, by leveraging the simplicity and accuracy of block-correlation channel modeling, we rigorously bound the performance limits of FBL-FAS from a statistical perspective, focusing on two key performance metrics: block error rate (BLER) and outage probability (OP).
-Furthermore, we introduce a novel complex-integral simplification method based on Gauss-Laguerre quadrature, which achieves higher approximation accuracy compared to existing Taylor-expansion-based approaches.
-Numerical results validate the robustness of the proposed analysis and clearly demonstrate the superiority of FBL-FAS over conventional multiple-antenna systems with fixed antenna placement.
+Finite blocklength (FBL) effects are often overlooked in analyses based on outage probability, where the assumption of vanishing decoding error can lead to overly optimistic performance predictions.
+Although fluid antenna systems (FASs) have been shown to deliver substantial spatial diversity for various system designs, their behavior under FBL constraints has not been adequately studied.
+Leveraging the simplicity, tractability, and accuracy of the spatial block-correlation channel model for FAS, this work investigates the random coding performance limits of FBL-coded systems equipped with fluid antennas.
+We analytically characterize how FAS contributes to performance enhancement under the FBL assumption and validate the findings through numerical evaluations.
+Both analytical and simulation results reveal that FAS offers pronounced advantages over conventional systems employing multiple fixed antennas in FBL-coded communication scenarios.
 I Introduction
-Fluid antenna systems (FAS) represent a communication architecture that achieves substantial spatial diversity gains by dynamically altering radiation characteristics through flexible antenna structures and adaptive protocol designs [1] .
-Unlike conventional approaches that seek to avoid antenna correlation in compact spaces, FAS intentionally embraces and exploits it to yield significant channel gains [2], thereby enabling robust service quality with reduced hardware overhead and simplified protocol design.
-Recent work [3] investigated the performance limits of FAS under the finite blocklength (FBL) regime, demonstrating that fluid antennas can simultaneously secure high spectrum and power efficiency.
-Furthermore, FBL theories [4, 5] provide both fundamental insights and practical guidelines for system design to meet the requirements of massive access, ultra-low latency, and high reliability.
-Unlike prior studies that primarily adopt outage probability (OP) as the performance metric for FAS, [3] introduces block error rate (BLER) as a universal and analytically tractable metric with a closed-form expression.
-The main scope and contributions of this work are summarized as follows:
-    In this work, we focus on the FBL-FAS problem [3] under the block-correlation channel model [6] of FAS.
-With accurate channel modeling, the performance limits of FBL-FAS are statistically characterized by providing tighter upper bounds without requiring prior channel state information (CSI).
-    We derive the PDF expression of the channel response after antenna/port selection, which is concisely represented by a non-central chi-square distribution.
-    Furthermore, we introduce a more general integral simplification method inspired by the Gauss-Laguerre approach.
-Compared with the previously used Taylor-expansion-assisted method [3], our method avoids amplified approximation errors from high-order expansion terms under large antenna correlation, and achieves improved approximation accuracy.
-The remainder of the paper is organized as follows: In Section II, FBL-FASs under block-correlation channel modeling are explained.
-Subsequently, our mains results are introduced in Section III while explicit derivations are deferred to Section VI.
-Numerical results are illustrated and explained in Section IV and the conclusion is drawn in Section V.
+Fluid antenna systems (FAS) represent a promising communication paradigm capable of achieving substantial spatial diversity by dynamically adjusting radiation characteristics through flexible antenna structures and adaptive protocol designs [1].
+Unlike traditional architectures that attempt to mitigate antenna correlation in compact spaces, FAS deliberately exploits such correlation to attain notable channel gains [2, 3], offering enhanced service reliability with lower hardware complexity and simplified signaling procedures.
+Most existing studies focus on outage probability performance [1, 4], which is not applicable when finite blocklength (FBL) effects are present [5, 6].
+Under the FBL regime, decoding errors do not vanish, and thus outage probability is not to be defined.
+Although [5] provides a closed-form block error rate (BLER) expression for random coding analysis, it requires extensive system-specific parameters, limiting its practical applicability.
+In this work, we investigate the random coding limits of FAS-enabled systems under FBL assumptions by carefully analyzing and simplifying both the BLER expression and the channel-strength distribution.
+The main contributions are summarized as follows:
+  • 1)
+    We derive the exact PDF of the channel strength $|g_{\mathrm{FAS}}|^{2}$ in expression (8), whose integral form is inherently complex.
+We also present a simplified numerical evaluation based on Gauss–Laguerre quadrature by (12).
+  • 2)
+    To further reduce complexity, we introduce a Gamma Approximation for the block-correlation model in (17), yielding a compact and computationally efficient PDF representation with high approximation accuracy.
+  • 3)
+    We derive a high-SNR approximation of the instantaneous BLER for short-packet/finite-length codes in (20), which eliminates the original integral form and substantially lowers computational overhead.
+The reproducible simulation code used in this work can be accessed at: https://github.com/BrooklynSEUPHD/Finite-blocklength-Fluid-Antenna-Systems-With-Spatial-Block-Correlation-Channel-Model.git.
+The rest of the paper is organized as follows.
+Sec.
+II introduces the block-correlation channel model for FBL-FAS.
+Sec.
+III presents the main analytical results with detailed proofs.
+Numerical evaluations are provided in Sec.
+IV, and conclusions are drawn in Sec.
+V.
 II System Descriptions
-In this work, we consider an uplink transmission scenario where a base station (BS) equipped with a conventional single antenna serves $U$ single-antenna user equipments with blocklength $M$.
-Each UE employs a fluid antenna with $N$ preset configurable locations/ports that are evenly distributed along a linear array of length $W$, where $W$ denotes the antenna length normalized by the wavelength at the operating frequency.
-Each port is regarded as an ideal antenna; however, due to limited hardware resources such as the radio frequency (RF) chain, only one port will be activated at a time.
-II-A Spatial Block-Correlation Channel Model for Single Antenna $N$-Ports FAS
-Our starting point builds upon the FBL-FAS analytical model in [3, 2].
-In contrast, we extend the analysis to incorporate the spatial block-correlation channel model in [6].
-Since this channel model differs from that in [3], it is necessary to illustrate how the block-correlation model can be adopted to generate random channel realizations within the FBL-FAS framework.
+In this work, we consider a point-to-point transmission scenario (downlink or uplink) where a base station (BS) equipped with a conventional single antenna serves a single-antenna user equipment for exchange of $U$ bits of information assuming transmission with blocklength $M$.
+The UE employs a fluid antenna with $N$ preset configurable locations/ports that are evenly distributed along a linear array of length $W$, where $W$ denotes the antenna length normalized by the wavelength at the operating frequency.
+Each port is regarded as an ideal antenna; however, due to limited hardware resources such as the radio frequency (RF) chain, only one port will be activated at a time, i.e., a single RF chain is assumed.
+II-A Finite-Blocklength Single Antenna FAS Signal Model
+The received signal can be expressed as
+$\boldsymbol{y}=g_{u,k}\boldsymbol{x}_{u}+\boldsymbol{\eta},$ (1)
+where $\boldsymbol{x}_{u}$ denotes the Gaussian codeword of the user with zero mean and variance $\sigma_{c}^{2}=\tfrac{1}{M}$, such that $\|\boldsymbol{x}_{u}\|^{2}_{2}=1$.
+The term $\boldsymbol{\eta}$ represents the additive white Gaussian noise (AWGN) with zero mean and variance $\tfrac{\sigma^{2}_{\eta}}{M}$, i.e., $\|\boldsymbol{\eta}\|^{2}_{2}=\sigma^{2}_{\eta}$.
+With a slight abuse of notation, $g_{u,k}$ denotes the $k$-th channel coefficient of the user, where $k\in\{1,2,\ldots,N\}$ represents the FAS port.
+The channel coefficients are Rayleigh distributed with variance $\sigma^{2}$ and correlated by correlation matrix $\boldsymbol{\Sigma}$.
+Moreover, the signal-to-noise ratio (SNR) is defined as:
+$\mathrm{SNR}=\frac{\mathbb{E}\left\{\|g_{u,k}\boldsymbol{x}_{u}\|^{2}_{2}\right\}}{\mathbb{E}\left\{\|\boldsymbol{\eta}\|^{2}_{2}\right\}}=\frac{\sigma^{2}}{\sigma^{2}_{\eta}}.$ (2)
+II-B Spatial Block-Correlation Channel Model for Single Antenna $N$-Ports FAS
+The spatial block-correlation channel model in [4] is considered for the potential simplicity, modeling accuracy and the tractability brought by the block-fitting.
+We first illustrate how the block-correlation model can be adopted to generate random channel realizations within the FBL-FAS framework.
 Let $\boldsymbol{g}_{u}\in\mathbb{C}^{N}$ denote the channel coefficient vector of the $u$-th user, and let $\boldsymbol{\Sigma}\in\mathbb{C}^{N\times N}$ represent the corresponding spatial correlation matrix.
-Owing to the even placement of ports, the correlation matrix $\boldsymbol{\Sigma}$ takes the form of a Toeplitz matrix:
-$\boldsymbol{\Sigma}=\begin{pmatrix}a(0)&a(1)&a(2)&...&a(N-1)\\ a(-1)&a(0)&a(1)&...&a(N-2)\\ \vdots&\ddots&&\vdots\\ a(-N+1)&a(-N+2)&...&a(-1)&a(0)\end{pmatrix}$, (1)
-where the generation function is:
-$a(n)=\operatorname{sinc}\left(\frac{2\pi nW}{N-1}\right).$ (2)
-With correlation matrix $\boldsymbol{\Sigma}$, one could generate $\boldsymbol{g}_{u}$ feasibly via eigenvalue-based construction [1, Eq. 5]:
+Owing to the even placement of ports and the assumption of Clarke’s correlation model, the matrix $\boldsymbol{\Sigma}$ takes the form of a Toeplitz matrix:
+$\small\boldsymbol{\Sigma}=\begin{pmatrix}a(0)&a(1)&a(2)&...&a(N-1)\\ a(-1)&a(0)&a(1)&...&a(N-2)\\ \vdots&\ddots&&\vdots\\ a(-N+1)&a(-N+2)&...&a(-1)&a(0)\end{pmatrix}$, (3)
+where the generation function is $a(n)=\operatorname{sinc}\left(\frac{2\pi nW}{N-1}\right)$.
+With correlation matrix $\boldsymbol{\Sigma}$, one could generate $\boldsymbol{g}_{u}$ feasibly via eigenvalue-based construction [1, Eq.
+5]:
 $\boldsymbol{g}_{u}=\boldsymbol{Q}\boldsymbol{\Lambda}^{\frac{1}{2}}\boldsymbol{g}_{0},$ (4)
-Here, $\boldsymbol{Q}$ denotes the matrix of left eigenvectors obtained from the eigenvalue decomposition of the correlation matrix $\boldsymbol{\Sigma}=\boldsymbol{Q}\boldsymbol{\Lambda}\boldsymbol{Q}^{\mathrm{H}}$, and $\boldsymbol{g}_{0}\in\mathbb{C}^{N}\sim\mathcal{CN}\left(\boldsymbol{0},\sigma^{2}\boldsymbol{I}\right)$.
+where $\boldsymbol{Q}$ denotes the matrix of left eigenvectors and $\boldsymbol{\Lambda}$ is the diagonal matrix of corresponding eigenvalues obtained obtained from the eigenvalue decomposition of the correlation matrix $\boldsymbol{\Sigma}=\boldsymbol{Q}\boldsymbol{\Lambda}\boldsymbol{Q}^{\mathrm{H}}$, and $\boldsymbol{g}_{0}\in\mathbb{C}^{N}\sim\mathcal{CN}\left(\boldsymbol{0},\sigma^{2}\boldsymbol{I}\right)$.
 However, directly analyzing the statistical distribution based on $\boldsymbol{\Sigma}$ is infeasible and overly complicated for any tractable analyses.
-Fortunately, the principal components of $\boldsymbol{\Sigma}$ can be captured by only a small number of dominant eigenvalues and their corresponding eigenvectors [6, Fig.
+Fortunately, the principal components of $\boldsymbol{\Sigma}$ can be captured by only a small number of dominant eigenvalues and their corresponding eigenvectors [4, Fig.
 3], [1].
 Hence, the correlation matrix $\boldsymbol{\Sigma}$ can be approximated by a block-diagonal matrix:
-$\widehat{\boldsymbol{\Sigma}}\in\mathbb{R}^{N\times N}=\begin{pmatrix}\mathbf{A}_{1}&\mathbf{0}&\cdots&\mathbf{0}\\ \mathbf{0}&\mathbf{A}_{2}&\cdots&\mathbf{0}\\ \vdots&&\ddots&\vdots\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{A}_{B}\end{pmatrix}$, (4)
+$\widehat{\boldsymbol{\Sigma}}\in\mathbb{R}^{N\times N}=\begin{pmatrix}\mathbf{A}_{1}&\mathbf{0}&\cdots&\mathbf{0}\\ \mathbf{0}&\mathbf{A}_{2}&\cdots&\mathbf{0}\\ \vdots&&\ddots&\vdots\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{A}_{B}\end{pmatrix}$.
+(5)
 Each diagonal block $\boldsymbol{A}_{b},~b\in\{1,2,\ldots,B\}$, is a constant correlation matrix of size $L_{b}\times L_{b}$, satisfying $\sum_{b=1}^{B}L_{b}=N$ (omit rounding errors).
 Intuitively, the correlation matrix characterizes the correlation among neighboring ports, where the correlation is typically strong due to the compact port spacing.
-The matrix $\widehat{\boldsymbol{\Sigma}}$ represents the fitted approximation of $\boldsymbol{\Sigma}$ according to [6, Eq.
+The matrix $\widehat{\boldsymbol{\Sigma}}$ represents the fitted approximation of $\boldsymbol{\Sigma}$ according to [4, Eq.
 20], while offering a more convenient structure for analytical treatment.
@@
 $\mathbf{A}_{b}=\begin{pmatrix}1&\mu_{b}^{2}&\cdots&\mu_{b}^{2}\\ \mu_{b}^{2}&1&\cdots&\mu_{b}^{2}\\ \vdots&&\ddots&\vdots\\ \mu_{b}^{2}&\cdots&\mu_{b}^{2}&1\end{pmatrix}$, (6)
-where $\mu_{b}^{2}$ is the correlation constant. Two important assumptions adopted throughout the remainder of the paper are summarized as follows:
-    Without loss of generality, $\mu_{b}^{2},~b\in\{1,2,\ldots\}$ are assumed identical, i.e., $\mu_{1}^{2}=\mu_{2}^{2}=\cdots=\mu^{2}\in\left(0.95,0.99\right)$, as established in [6, Sec. III-B-2)].
-    The block size $L_{b}$ is adaptive to the desired approximation accuracy, which follows the procedure in [6, Algorithm 1].
-For implementation details, the reader is kindly referred to the open-source code provided therein.
-II-B Finite-Blocklength Single Antenna FAS Signal Model
-Assuming synchronous transmission, $U$ single-antenna users transmit $M$-length signals, and the received signal can be expressed as
-$\boldsymbol{y}=g_{i,k}\boldsymbol{x}_{i}+\sum_{u\neq i}^{U}g_{u,k}\boldsymbol{x}_{u}+\boldsymbol{\eta},$ (6)
-where $\boldsymbol{x}_{u}$ denotes the Gaussian codeword of user $u$ with zero mean and variance $\sigma_{c}^{2}=\tfrac{1}{M}$, such that $\|\boldsymbol{x}_{u}\|^{2}_{2}=1$.
-The term $\boldsymbol{\eta}$ represents the additive white Gaussian noise (AWGN) with zero mean and variance $\tfrac{\sigma^{2}_{\eta}}{M}$, i.e., $\|\boldsymbol{\eta}\|^{2}_{2}=\sigma^{2}_{\eta}$.
-With a slight abuse of notation, $g_{u,k}$ denotes the $k$-th channel coefficient of user $u$, where $k\in\{1,2,\ldots,N\}$.
-The channel coefficients are Rayleigh distributed with variance $\sigma^{2}$ and correlated by (4).
-Moreover, the signal-to-noise ratio (SNR) is defined as
-$\mathrm{SNR}=\frac{\mathbb{E}\left\{\|g_{i,k}\boldsymbol{x}_{i}\|^{2}_{2}\right\}}{\mathbb{E}\left\{\|\boldsymbol{\eta}\|^{2}_{2}\right\}}=\frac{\sigma^{2}}{\sigma^{2}_{\eta}}.$ (7)
+where $\mu_{b}^{2}\in\left(0.95,0.99\right),~b\in\{1,2,\ldots\}$ is the correlation constant.
+In summary, in this work, the Clarke’s correlation model in (3) is adopted and the corresponding block diagonal approximation in (5) is adopted for theoretical tractability.
+III Main Results
+This section introduces our main findings in terms of PDF derivations and fast calculation in Sec.
+III-A, further simplified expressions for ease of computation in Sec.
+III-B and BLER bound approximation in Sec.
+III-B.
+The goal is to achieve simplified statistical results while maintaining high prediction accuracy.
+III-A PDF of $|g_{\mathrm{FAS}}|^{2}$ and BLER under Block-Correlation Model
+The PDF of the random variable $|g_{\mathrm{FAS}}|^{2}=\max\left\{|g_{u,1}|^{2},|g_{u,2}|^{2},\ldots,|g_{u,N}|^{2}\right\}$ for FBL-FAS under the block-correlation channel model is given in (8), where $f_{\chi_{2}^{\prime 2}}(x;\lambda)$ and $F_{\chi_{2}^{\prime 2}}(x;\lambda)$ denote the PDF and CDF of a non-central chi-square distribution with $2$ degrees of freedom and non-centrality parameter $\lambda$, respectively.
+We explain the derivations leading to (8) at the top of next page.
+Specifically, we derive the PDF from the CDF expression given next.
+Remark: Readily converted from [4, Eq. 43], [9], the CDF of random variable $|g_{\mathrm{FAS}}|^{2}=\max\left\{|g_{u,1}|^{2},|g_{u,2}|^{2},\ldots,|g_{u,N}|^{2}\right\}$ is:
+$\begin{aligned} &F_{|g_{\mathrm{FAS}}|^{2}}=P\left\{|g_{\mathrm{FAS}}|^{2}\leq t\right\}\\ &=\prod_{b=1}^{B}\int_{0}^{\infty}\frac{e^{-\frac{r_{b}}{2}}}{2}\times\left[1-Q_{1}\left(\sqrt{\frac{\mu^{2}r_{b}}{1-\mu^{2}}},\sqrt{\frac{t}{1-\mu^{2}}}\right)\right]^{L_{b}}\mathrm{d}r_{b}\end{aligned}.$ (7)
 $f_{|g_{\mathrm{FAS}}|^{2}}(t)$ (8)
 $=\sum_{b=1}^{B}\left\{\int_{0}^{\infty}\frac{e^{-\frac{r_{b}}{2}}L_{b}}{2}\left[F_{\chi_{2}^{\prime 2}}(\frac{t}{1-\mu^{2}};\frac{\mu^{2}r_{b}}{1-\mu^{2}})\right]^{L_{b}-1}\frac{f_{\chi_{2}^{\prime 2}}(\frac{t}{1-\mu^{2}};\frac{\mu^{2}r_{b}}{1-\mu^{2}})}{1-\mu^{2}}\mathrm{d}r_{b}\right\}\times\prod_{j\neq b}^{B}\left\{\int_{0}^{\infty}\frac{e^{-\frac{r_{j}}{2}}}{2}\left[F_{\chi_{2}^{\prime 2}}\left(\frac{t}{1-\mu^{2}};\frac{\mu^{2}r_{j}}{1-\mu^{2}}\right)\right]^{L_{j}}\mathrm{d}r_{j}\right\},$
-III Main Results
-In this section, we derive the PDF of the random variable $|g_{\mathrm{FAS}}|^{2}=\max\left\{|g_{u,1}|^{2},|g_{u,2}|^{2},\ldots,|g_{u,N}|^{2}\right\}$, based on which both the conditional BLER (with known channel response) and the statistical BLER (without prior CSI) are obtained.
-Furthermore, we investigate the outage OP expressions of FBL-FAS under the block-correlation channel model.
-In addition, we clarify why the integral simplification method in [3, Eq.
-52], i.e., the Taylor-expansion-assisted mean value theorem for integrals, is not applicable to the block-correlation model.
-Instead, we propose a tailored simplification approach for block correlation model based on Gauss–Laguerre quadrature.
-Figure 1: Illustration of block-correlation model’s CDF and the PDF (exact CDF in (13), derived PDF in (8) and Monte-Carlo samples.) with $N=10,~\mu^{2}=0.97,~W=0.5$.
-III-A PDF of $|g_{\mathrm{FAS}}|^{2}$ and BLER under Block-Correlation Model
-Result 1: The PDF of the random variable $|g_{\mathrm{FAS}}|^{2}=\max\left\{|g_{u,1}|^{2},|g_{u,2}|^{2},\ldots,|g_{u,N}|^{2}\right\}$ for FBL-FAS under the block-correlation channel model is given in (8), where $f_{\chi_{2}^{\prime 2}}(x;\lambda)$ and $F_{\chi_{2}^{\prime 2}}(x;\lambda)$ denote the PDF and CDF of a non-central chi-square distribution with $2$ degrees of freedom and non-centrality parameter $\lambda$, respectively.
-Proof: See Appendix VI-A. $\blacksquare$
-For illustration, the exact CDF in (13), the derived PDF in (8), and the Monte Carlo samples obtained via (3) are plotted in Fig.
-1, where the channel model setups are generated with $N=10$, $\mu^{2}=0.97$, and $W=0.5$.
-Notably, consistent with [6], the CDF in (13) assumes $\sigma^{2}=2$, which will be adopted throughout the simulations in this work [^1] It is straightforward to extend the CDF and PDF to different variances $\sigma^{2}$, which can be done by a standard change of variables in the PDF/CDF..
-Remark 1 According to [3, Theorem 5], for FBL-FAS with $U$ codewords to be detected under transmision of $M$ blocklength, the BLER can be upperbounded by the conditional (known channel response) error probability denoted by $\epsilon$:
-$\epsilon\leq\sum_{{U}^{\prime}=0}^{U}\frac{{U}^{\prime}}{U}e^{{L}^{\prime}-M\log\left(1+\frac{0.25\sigma_{\eta^{\prime}}^{2}}{\sigma_{\eta}^{2}}\right)},$ (9)
-where ${L}^{\prime}=\log\binom{U}{{U}^{\prime}}^{2}=\sum_{i=1}^{{U}^{\prime}-1}\log\frac{U-i}{{U^{\prime}-i}}$ and $\sigma^{2}_{\eta^{\prime}}=2{U}^{\prime}\sigma_{c}^{2}|g_{\mathrm{FAS}}|^{2}$.
-With the conditional BLER in (9), one could easily obtain the statistical BLER (without CSI) denoted by $\bar{\epsilon}$ below:
-$\bar{\epsilon}=\sum_{{U}^{\prime}=0}^{U}\frac{{U}^{\prime}}{U}\int_{0}^{\infty}f_{|g_{\mathrm{FAS}}|^{2}}(t)e^{{L}^{\prime}-M\log\left(1+\frac{0.25\sigma_{\eta^{\prime}}^{2}}{\sigma_{\eta}^{2}(t)}\right)}\mathrm{d}t,$ (10)
-where $f_{|g_{\mathrm{FAS}}|^{2}}(t)$ is the PDF of $|g_{\mathrm{FAS}}|^{2}$ and $\sigma_{\eta}^{2}(t)=2{U}^{\prime}\sigma_{c}^{2}t$.
-III-B SINR-based Outage Probability and Integral Simplification
-Remark 2 Let $\Gamma$ denote the SINR under FBL-FAS and $\gamma_{th}$ as the OP threshold, in accordance with [3, Appendix H], and let the OP event be defined as::
-$\left\{\Gamma\leq\gamma_{\mathrm{th}}\right\}$ (11)
-$=\left\{|g_{\mathrm{FAS}}|^{2}\leq\frac{\sigma_{\eta}^{2}\gamma_{\mathrm{th}}}{1-\left(U-1\right)\left(U\bar{\rho}+1\right)\gamma_{\mathrm{th}}}\right\},$
-where $\bar{\rho}=\sqrt{\frac{\pi}{4M}}$ is the average codeword correlation (please refer to [3, Theorem 3] for details).
-Hereby, with CDF in (13), the OP for FBL-FAS under block-correlation channel model is calculated as:
-$p^{\mathrm{FBL-FAS}}_{\mathrm{out}}$ (12)
-$=\prod_{b=1}^{B}\int_{0}^{\infty}\frac{e^{-\frac{r_{b}}{2}}}{2}\times\left[1-Q_{1}\left(\sqrt{\frac{\mu^{2}r_{b}}{1-\mu^{2}}},\sqrt{\frac{t_{th}}{1-\mu^{2}}}\right)\right]^{L_{b}}\mathrm{d}r_{b},$
-where $t_{th}=\frac{\sigma_{\eta}^{2}\gamma_{\mathrm{th}}}{1-\left(U-1\right)\left(U\bar{\rho}+1\right)\gamma_{\mathrm{th}}}$.
-Still, the expression in (12) has an infeasible product-integral structure.
-We next find an effective method to simplify the calculation.
-Result 2: For the integral in (12), the simplification method originally adopted in [3, Theorem 7]—namely, the Taylor-expansion-assisted mean value theorem for integrals—is not applicable to the block-correlation model.
-In this work, we propose a novel approach for integral simplification based on Gauss–Laguerre quadrature, tailored to the analysis of FBL-FAS under the block-correlation model.
-Details: See Appendix VI-B. $\blacksquare$
-Figure 2: Illustrations of BLER performance:
--Different Lines: Red solid lines denote the statistical BLER of FBL-FAS, red markers represent the corresponding conditional BLER, and black dotted lines indicate the conditional BLER of the conventional $L$-antenna system.
--Different Figures: 1) BLER vs $U$: $W=1$, $M=5$, $N\in\{5,50\}$, $L\in\{1,2\}$ , and $\mathrm{SNR}=20$ dB; 2) BLER vs SNR: $W=0.5$, $M=5$, $U=10$, $N\in\{5,50,1000\}$, and $L\in\{1,2\}$; 3) BLER vs $N$: $W=1$, $M=5$, $U=10$, $L\in\{1,2\}$, and $\mathrm{SNR}=12$ dB.
-4) BLER vs $W$: $M=5$, $U=10$, $N=5000$, $L\in\{1,2\}$, and $\mathrm{SNR}=12$ dB.
-Figure 3: Illustration of OP performance:
--Different Lines: Red solid lines denote the OP of FBL-FAS, while black dotted lines indicate the OP of the MRC system.
--Different Figures: Up: OP versus SNR (dB) with $W=0.5$, $N\in\{5,50,500\}$, $L\in\{1,3,5\}$, $U=20$, $M=5$, and $\gamma_{th}=10^{-3}$; Down: OP versus the number of codewords $U$ with $W=0.5$, $N\in\{5,50,500,1000\}$, $L\in\{1,3,5\}$ (conventional antenna system), $\mathrm{SNR}=-35$ dB, $M=5$, and $\gamma_{th}=10^{-4}$.
-IV Numerical Results
-In this part, we illustrate the BLER and OP performance of FBL-FAS under the block-correlation model.
-The OP and BLER for a conventional $L$-antenna scheme using maximum ratio combining (MRC) [3, Section III-D] are included as benchmarks.
-For the Monte Carlo simulations, random channel realizations are generated according to (3).
-Unless otherwise specified, the default setup assumes channel variance $\sigma^{2}=2$ (consistent with [6]) and correlation parameter $\mu^{2}=0.97$.
-IV-A BLER and OP Performance
-The BLER and OP performance results are illustrated in Fig. 2 and Fig. 3, respectively, with parameter setups and explanations on curve and marker types in the caption part:
-In Fig.
-2, BLER results are presented under different numbers of codewords $U$, SNR levels, and numbers of ports $N$.
-Several common findings can be summarized: 1) The BLER can be significantly reduced with denser port placement and wider antenna length, i.e., increasing $N$ or $W$ improves the BLER performance, which can easily surpass that of conventional single-antenna systems and even multi-antenna systems.
-2) With a sufficiently large number of ports $N$, the block-correlation channel model provides accurate predictions of BLER performance through the statistical expression in (10), thereby validating both the proposed approach and the accuracy of the block-correlation channel model.
-In Fig.
-3, the OP performance is illustrated under different SNR values and numbers of active codewords $U$.
-The key observations are as follows: 1) The single-antenna MRC system experiences an error floor due to user interference in finite-blocklength transmission.
-2) FBL-FAS exhibits a much steeper OP decay as the SNR increases.
-Notably, even single-antenna FBL-FAS can achieve lower OP than a 5-antenna MRC system when $50$ or $500$ ports are available.
-3) Compared with the 1-antenna MRC system, FBL-FAS achieves $10^{7}$ to $10^{4}$ times lower OP, and it even surpasses the 3-antenna MRC system with approximately 10 times better OP performance.
-IV-B Integral Approximtion
-To investigate the impact of $\mu$ on the approximation accuracy of different methods, we examine the integral component
-$\int_{0}^{\infty}\frac{e^{-\frac{r_{b}}{2}}}{2}\left[1-Q_{1}\!\left(\sqrt{\tfrac{\mu^{2}r_{b}}{1-\mu^{2}}},\sqrt{\tfrac{t}{1-\mu^{2}}}\right)\right]^{L_{b}}\mathrm{d}r_{b}$
-from (13) with $L_{b}=3$ and quadrature expansion order of $32$, as illustrated in Fig.
-4.
-It can be observed that the Taylor expansion-assisted method becomes inaccurate as $\mu$ increases from $0.1$ to $0.97$.
-In contrast, the proposed Gauss–Laguerre-based approach remains effective across the entire range of correlation parameters.
-Figure 4: Illustration of integral approximation under different correlation values in the block-correlation model, using the Taylor expansion-assisted method [3, Theorem 7] and the proposed Gauss–Laguerre method.
-The figures in the first row show the integral approximation while the figures in the second row show the corresponding approximation error, $|\tilde{f}(t)-f(t)|^{2}$.
-The $x$-axis denotes the range of $t$.
-V Conclusion
-In this work, the performance limits of FBL-FAS [3] are tightly bounded from a statistical perspective by considering the block-correlation channel model in [6].
-The PDF expression of the channel gain after port selection is derived and concisely characterized by a non-central chi-square distribution, whose evaluation is further simplified using the Gauss-Laguerre approach to yield accurate approximations.
-Numerical results validate the tightness of the statistical bounds and highlight the significant spectrum and energy efficiency achieved by FASs in the FBL regime.
-Compared with conventional $L$-antenna systems, FBL-FAS emerges as a promising enabler for future wireless networks.
-VI Appendix
-VI-A Proof of Results 1
-Remark A-1: Readily converted from [6, Eq. 43], the CDF of random variable $|g_{\mathrm{FAS}}|^{2}=\max\left\{|g_{u,1}|^{2},|g_{u,2}|^{2},\ldots,|g_{u,N}|^{2}\right\}$ is:
-$\begin{aligned} &F_{|g_{\mathrm{FAS}}|^{2}}=P\left\{|g_{\mathrm{FAS}}|^{2}\leq t\right\}\\ &=\prod_{b=1}^{B}\int_{0}^{\infty}\frac{e^{-\frac{r_{b}}{2}}}{2}\times\left[1-Q_{1}\left(\sqrt{\frac{\mu^{2}r_{b}}{1-\mu^{2}}},\sqrt{\frac{t}{1-\mu^{2}}}\right)\right]^{L_{b}}\mathrm{d}r_{b}\end{aligned}.$ (13)
-Subsequently, we derive the corresponding PDF by Remark A-1. For ease of descriptions, we denote (13) by:
+Figure 1: Illustration of block-correlation model’s CDF and the PDF (block diagonal CDF in (7), the derived PDF in (8) and the corresponding Monte-Carlo samples.) with $N=10,~\mu^{2}=0.97,~W=0.5$.
+For ease of descriptions, we denote (7) by:
 $F_{|g_{\mathrm{FAS}}|}=\prod_{b=1}^{B}G_{b}(t),$ (9a)
@@
     where $f_{\chi_{2}^{\prime 2}}(x;a)=\frac{1}{2}e^{-\frac{x+a}{2}}I_{0}(\sqrt{ax})$ is the corresponding chi-square distribution’s PDF.
-Thereby, if one consider indentity of (15) on (14a) and then use the derivative rule $f_{|g_{\mathrm{FAS}}|^{2}}(t)\equiv\frac{\mathrm{d}F_{|g_{\mathrm{FAS}}|^{2}}(t)}{\mathrm{d}t}=\sum_{b=1}^{B}\left(\frac{\mathrm{d}G_{b}(t)}{\mathrm{d}t}\right)\prod_{\begin{array}[]{c}j\neq b\end{array}}^{B}G_{j}(t)$ with identity in (16), (8) is obtained, which completes the proof on Result 1 in Section III-A.
-VI-B Explanations on Result 2
-Different from the constant channel correlation model in [3], the block-approximation model assumes the channel correlation to be a constant close to $1$, which aims to capture the correlation among ports in close vicinity.
-However, when $\mu^{2}\rightarrow 1$, if one continues to adopt the “frozen” point integral simplification method in [3, Theorem 7], the approximation becomes inaccurate due to the limitations of the Taylor expansion in [3, Eq.
-(56)] for $\mu^{2}\rightarrow 1$.
-Specifically, for the function component $1-Q_{1}\!\left(\sqrt{\tfrac{\mu^{2}r_{b}}{1-\mu^{2}}},\sqrt{\tfrac{t}{1-\mu^{2}}}\right)$ the constant terms $\tfrac{\mu^{2}}{1-\mu^{2}}\rightarrow+\infty$ or $\tfrac{1}{1-\mu^{2}}\rightarrow+\infty$ will excessively amplify the higher-order error terms after Taylor expansion, leading to failure of the approximation.
-Therefore, it is both useful and interesting to explore ways to compensate for the drawbacks of Taylor-expansion-based methods.
-In this regard, we introduce how to simplify the integral calculation of expressions with structures similar to (12), inspired by the Gauss–Laguerre quadrature method [7, 8]:
-    For an integrand of the form $e^{-x}f(x)$, the integral can be approximated as
-    $\int_{0}^{\infty}e^{-x}f(x)\,\mathrm{d}x\;\approx\;\sum_{i=1}^{N_{GL}}w_{i}f(x_{i}),$ (17)
+Thereby, if one consider indentity (10) in (9a) and then use the derivative rule $f_{|g_{\mathrm{FAS}}|^{2}}(t)\equiv\frac{\mathrm{d}F_{|g_{\mathrm{FAS}}|^{2}}(t)}{\mathrm{d}t}=\sum_{b=1}^{B}\left(\frac{\mathrm{d}G_{b}(t)}{\mathrm{d}t}\right)\prod_{\begin{array}[]{c}j\neq b\end{array}}^{B}G_{j}(t)$ with identity in (11), (8) is obtained, which completes the proof.
+While (8) manifests complex integral structures, it can be feasibly computed by the Gauss–Laguerre quadrature method [7, 8] as follows:
+    For an integrand of the form $e^{-x}f(x)$, the integral can be approximated as:
+    $\int_{0}^{\infty}e^{-x}f(x)\,\mathrm{d}x\;\approx\;\sum_{i=1}^{N_{\mathrm{GL}}}w_{i}f(x_{i}),$ (12)
     where $x_{i}$ are the nodes and $w_{i}$ are the corresponding weights.
-The parameter $N_{GL}$ is the expansion order, which determines the approximation accuracy (typically, choosing $N_{GL}$ up to $64$ is sufficient).
+The parameter $N_{\mathrm{GL}}$ is the expansion order, which determines the approximation accuracy (typically, choosing $N_{\mathrm{GL}}$ up to $64$ is sufficient but it can be well-tuned catering to different requirements).
     According to the Golub–Welsch theorem [7, 8], the values of $(x_{i},w_{i})$ can be derived by performing an eigenvalue decomposition on the following Jacobi matrix:
@@
 The corresponding weights $w_{i}$ are obtained from the squared elements of the first row of $\boldsymbol{W}$.
+For illustration of block-correlation modeling, the CDF in (7), the derived PDF in (8), and the Monte Carlo samples (generated via block-correlation approximation) obtained from (4) are plotted in Fig.
+1, where the channel model setups are generated with $N=10$, $\mu^{2}=0.97$, and $W=0.5$.
+Notably, consistent with [4], the CDF in (7) assumes $\sigma^{2}=2$, which is adopted throughout the simulations in this work [^1] It is straightforward to extend the CDF and PDF to different variances $\sigma^{2}$, which can be done by a standard change of variables in the PDF/CDF..
+III-B Simplified PDF of $|g_{\mathrm{FAS}}|^{2}$ via Gamma Approximation
+As observed in Fig.
+1, for $t\geq 0$ the distribution of $f_{|g_{\mathrm{FAS}}|^{2}}(t)$ is unimodal and right-skewed, a key characteristic of the Gamma distribution family.
+This motivates the use of a Gamma approximation to simplify (8) and substantially reduce computational complexity.
+To approximate the random variable $t=|g_{\mathrm{FAS}}|^{2}$ by a Gamma distribution, we first compute its expectation and variance:
+$\mathbb{E}\{t\}$ $=\int_{0}^{\infty}t\,f_{|g_{\mathrm{FAS}}|^{2}}(t)\,\mathrm{d}t,$ (14)
+$\mathrm{Var}\{t\}$ $=\mathbb{E}\{t^{2}\}-\left(\mathbb{E}\{t\}\right)^{2}$
+$=\int_{0}^{\infty}t^{2}f_{|g_{\mathrm{FAS}}|^{2}}(t)\,\mathrm{d}t-\left(\int_{0}^{\infty}t\,f_{|g_{\mathrm{FAS}}|^{2}}(t)\,\mathrm{d}t\right)^{2}.$
+We then assume $t\sim\mathcal{G}(k,\theta)$ with PDF and CDF:
+$f_{\mathcal{G}}(t)$ $=\frac{1}{\Gamma(k)\theta^{k}}t^{k-1}e^{-t/\theta},$ (15)
+$F_{\mathcal{G}}(t)$ $=\frac{\gamma\!\left(k,t/\theta\right)}{\Gamma(k)},\quad t\geq 0,$
+where $\Gamma(\cdot)$ and $\gamma(\cdot,\cdot)$ denote the Gamma function and lower incomplete Gamma function, respectively.
+Matching the first two moments of the Gamma distribution with those in (14) yields:
+$k\theta$ $=\mathbb{E}\{t\},\qquad k\theta^{2}=\mathrm{Var}\{t\},$ (16)
+$\Rightarrow\quad k$ $=\frac{\mathbb{E}\{t\}^{2}}{\mathrm{Var}\{t\}},\qquad\theta=\frac{\mathrm{Var}\{t\}}{\mathbb{E}\{t\}}.$
+With these parameters, the PDF and CDF of the channel strength in (8) and (7) can be approximated as:
+$f_{|g_{\mathrm{FAS}}|^{2}}(t)\approx\frac{1}{\Gamma(k)\theta^{k}}t^{k-1}e^{-t/\theta},F_{|g_{\mathrm{FAS}}|^{2}}(t)\approx\frac{\gamma\!\left(k,t/\theta\right)}{\Gamma(k)},$ (17)
+where $k$ and $\theta$ are determined by (16).
+To illustrate the accuracy of the proposed approximation, Fig.
+2 compares the exact CDF/PDF, the Gauss–Laguerre approximation with $N_{\mathrm{GL}}=128$ weights from (12), and the Gamma-approximated PDF in (17), for $N=10$, $\mu^{2}=0.97$, and $W=0.5$.
+The results confirm both the high fidelity of the Gauss–Laguerre approach and the effectiveness of the proposed Gamma-approximation method.
+Finally, we remark on computational complexity.
+Computing the exact PDF via Gauss–Laguerre quadrature requires $\mathcal{O}(B\,N_{\mathrm{GL}})$ operations per evaluation point $t$.
+In contrast, the proposed Gamma-approximated PDF is in closed form and requires only $\mathcal{O}(1)$ operations, with a negligible one-time cost for moment matching.
+This makes the Gamma approximation highly attractive for large-scale or real-time evaluations.
+Figure 2: Illustration of block-correlation model’s CDF and the PDF (exact CDF in (7), exact PDF in (8), the Gauss-Laguerre-approximated PDF with $N_{\mathrm{GL}}=128$ by (12) and the Gamma-distribution-approximated PDF by (17).) with $N=10,~\mu^{2}=0.97,~W=0.5$.
+III-C BLER Approximation
+Remark: For a given instantaneous SNR $\gamma$, the finite-blocklength BLER can be expressed as [5]:
+$\epsilon(\gamma)=Q\!\left(\frac{C(\gamma)-R_{c}}{\sqrt{V_{\mathrm{dis}}(\gamma)/M}}\right),$ (18)
+where $C(\gamma)=\frac{1}{2}\log_{2}(1+\gamma)$ is the channel capacity, $V_{\mathrm{dis}}(\gamma)=\frac{\gamma}{2}\frac{\gamma+2}{(\gamma+1)^{2}}\log_{2}^{2}(e)$ is the channel dispersion, $R_{c}=\frac{U}{M}$ is the code rate, and $Q(x)=\int_{x}^{\infty}\frac{1}{\sqrt{2\pi}}e^{-\tfrac{t^{2}}{2}}\mathrm{d}t$ is the standard $\mathrm{Q}$-function.
+We first rewrite the instantaneous SNR for fluid antennas.
+Since port selection yields the channel strength $t=|g_{\mathrm{FAS}}|^{2}$, the instantaneous SNR becomes $\gamma(t)=\frac{|g_{\mathrm{FAS}}|^{2}}{\sigma^{2}_{\eta}}$, where the PDF and CDF of $t$ have been approximated in (17).
+To avoid the computational burden associated with the $Q$-function, we derive a high-SNR approximation of (18). Using $\gamma\to\infty$ and the Gaussian tail approximation, we obtain:
+$V_{\mathrm{dis}}(\gamma)$ $=\frac{\gamma}{2}\frac{\gamma+2}{(\gamma+1)^{2}}\log_{2}^{2}(e)~\xrightarrow[\gamma\to\infty]{}~V_{\infty}\triangleq\frac{1}{2}\log_{2}^{2}(e),$ (19)
+$Q(x)$ $\approx\frac{1}{\sqrt{2\pi}\,x}\exp\!\left(-\frac{x^{2}}{2}\right),\qquad x\to+\infty.$
+Therefore, the high-SNR BLER approximation is:
+$\epsilon(\gamma)\approx\frac{1}{\sqrt{2\pi}\,x(\gamma)}\,\exp\!\left(-\frac{x^{2}(\gamma)}{2}\right),$ (20)
+where $x(\gamma)=\sqrt{\frac{M}{V_{\infty}}}\Biggl(\frac{1}{2}\log_{2}(1+\gamma)-R_{c}\Biggr)$ and $V_{\infty}=\frac{1}{2}\log_{2}^{2}(e)$.
+This simplified representation eliminates the integral in (18) and greatly reduces computational complexity.
+Notably, approximation in (19) is expected to scale up the BLER and upperbounds the exact one.
+Figure 3: BLER comparison between exact and high-SNR approximation for $\sigma^{2}=2$, $|g_{\mathrm{FAS}}|^{2}=1$, $U=100$, and $M=400$.
+Fig.
+3 compares the exact BLER from (18) and the approximation in (20).
+As expected, the high-SNR approximation is simplified but manifests higher BLER than the exact expression since $\frac{\gamma}{2}\frac{\gamma+2}{(\gamma+1)^{2}}\rightarrow 1$ only when $\gamma\rightarrow+\infty$.
+Since both (18) and (20) depend on the channel strength $|g_{\mathrm{FAS}}|^{2}$, whose PDFs (exact and approximated) are given in (8) and (17), the average BLER can now be efficiently evaluated as:
+$\bar{\epsilon}=\int_{0}^{+\infty}f_{|g_{\mathrm{FAS}}|^{2}}(t)\,\epsilon(t)\mathrm{d}t,$ (21)
+where all required components have been significantly simplified, enabling fast and accurate BLER prediction for fluid-antenna-enabled systems.
+IV Numerical Results
+In this section, simulations are conducted to demonstrate the advantages of FAS under the FBL assumption.
+The spatial diversity gain is verified by comparison with a conventional $L$-antenna system (fixed and independent antennas).
+The BLER of the conventional $L$-antenna system can be evaluated by substituting $\gamma=\frac{t_{L}}{\sigma^{2}_{\eta}}$ into (18), where $t_{L}=\sum_{l=1}^{L}|g_{l}|^{2}$ denotes the channel strength after equalization (e.g., maximum-ratio combining/transmission) and $g_{l},l\in\{1,\ldots,L\}$ are independent Rayleigh variables with zero mean and variance $\sigma^{2}=2$.
+Accordingly, $t_{L}\sim\mathcal{G}(L,\sigma^{2})$ follows a Gamma distribution with PDF $f_{t_{L}}(t)=\frac{1}{\Gamma(L)\sigma^{2L}}t^{L-1}e^{-t/\sigma^{2}},~t\geq 0$.
+The average BLER of the conventional $L$-antenna system is:
+$\mathrm{Benchmark:}~\bar{\epsilon}_{L}=\int_{0}^{+\infty}\epsilon_{t_{L}}(t)f_{t_{L}}(t)\mathrm{d}t.$ (22)
+Notably, FBL-FAS activates only a single antenna (one RF chain), whereas a conventional system is equiped with $L$ RF chains whose advantage is conspicuous compared to FAS.
+Unless otherwise specified, $U=100$ and $L=400$ are fixed, yielding a code rate of $\frac{100}{400}=0.25$ for a stringent short-packet scenario.
+The block-correlation constant is set as $\mu^{2}=0.97$.
+Figure 4: BLER versus SNR with $N=10$ available ports, $L\in\{1,3,5\}$, and $W=1$.
+All simulations are conducted within block-correlation model and the exact PDF refers to block-diagonal expression in (8) and the approximated PDF refers to the Gamma-approximated expression in (17).
+Fig.
+4 compares FBL-FAS and conventional systems with $L\in\{1,3,5\}$ when $N=10$ and $W=1$.
+The accuracy of both the channel-strength PDF approximation and the BLER expression is confirmed.
+Moreover, FBL-FAS exhibits significant random coding gain over the conventional multi-antenna system.
+Figure 5: BLER versus aperture with $N=50$ available ports, $L\in\{1,3,5,10,20\}$, and SNR =10=10 dB.
+Fig.
+5 shows the BLER of FBL-FAS versus aperture with $N=50$ and SNR =10=10 dB.
+For $W=5$, the minimum spacing between ports is $0.1$ (normalized) under uniform placement.
+Notably, with $W=5$, the FBL-FAS achieves performance comparable to a conventional system with $20$ antennas.
+With increasing aperture and fixed $N$, the available spatial diversity grows and the BLER decreases monotonically.
+It is also observed that the PDF approximation error increases when comparing Fig.
+4 and Fig.
+5.
+When $N=10$, the gap between the exact and approximate PDFs is minor, whereas it becomes larger for $N=50$.
+This is expected since the Gamma approximation in (17) is implicitely correlated to $N$ and the envolop of Gamma distribution is relatively regimented while (8) becomes increasingly unimodal with larger $N$.
+In practice, $N$ is constrained by the limit of hardware design, and moderate configurations (e.g., Fig.
+4) provide accurate performance prediction with reduced complexity.
+V Conclusion
+This work investigates the random coding performance limit of FBL communication with fluid antennas.
+Both the channel-strength model and a tractable BLER expression are developed to enable efficient performance prediction.
+We derive the PDF of block-correlated FAS and present an accurate approximation.
+In addition, the BLER expression is simplified to an integral-free form, significantly reducing computational complexity.
+Numerical results confirm that the spatial diversity provided by fluid antennas yields substantial gains for short-packet communication systems, highlighting a promising design paradigm.